find the measure of each interior angle of a regular 15-sided poygon
Answer by jim_thompson5910(35256)

I = (180(n-2))/n

I = (180(15-2))/(15)

I = (180(13))/(15)

I = 2340/(15)

I = 156

So the measure of each interior angle of a 15-sided polygon is 156 degrees.